# Towards Enforcing Company Policy Adherence in Agentic Workflows

### v1

@@ 1 Introduction
Our approach enforces policy adherence through a two-phase process illustrated in Figure 1: (1) an offline buildtime step automatically maps policy fragments to the relevant tools and generates policy validation code; (2) during runtime these validators—called ToolGuards—are executed before agent’s tool invocation. If a planned action violates a policy, the agent is prompted to self-reflect and revise its plan before proceeding. Ultimately, the deployed ToolGuards will prevent the agent from taking an action violating a policy.
We evaluate the proposed approach on the $\tau$-bench Airlines domain (Yao et al., 2024) with a large set of (naturalistic, real-world) policies, 14 tools, and diverse customer-service scenarios. [^2] The other $\tau$-bench domain of Retail is smaller and less diverse; hence we limit the scope of this work to Airlines. We show that this ambitious and challenging task pushes the limits of contemporary LLMs, present encouraging preliminary results, and outline directions for future research and practical steps needed for this approach to mature toward enterprise adoption. Our code and data will be made available.
@@ 2 Related Work
Prior art in the area of constraint and policy adherence in agentic workflows is very sparse, with two major studies, to the best of our knowledge, explicitly addressing the integration of company policies within agentic flows (contrary to post-hoc remedy). Yao et al. (2024) were among the pioneers in this field, introducing a framework for evaluation of company policies adherence in multi-turn simulated flows in two domains: Airlines and Retail. Policies are represented by short documents, where domain-specific rules and constraints are detailed in natural language. Although the study reports failure rates stemming from a range of issues, the authors estimate that 25% of failures occur since the agent "fails to understand the domain-specific knowledge or rules". A recent work by Li et al. (2025) offers a framework for evaluation of policy adherence in multi-step single-turn tasks in five diverse domains. The authors report relatively low pass-rate across agents, varying from 31% to 69%. Both studies introduce policies to the agent in its prompt, instructing the LLM to follow the guidelines while satisfying user requirements.
Contrary to the pro-active policy enforcement above, another line of studies take the post-hoc reflection approach, by designing guardrails that check if any action taken by an agent violates a predefined policy (Huang et al., 2024; Rothkopf et al., 2024; Hua et al., 2024). While this strategy is often easier for implementation (e.g., detecting that an unauthorized refund was issued is simpler than preventing it), it may fall short of the strict demands of enterprise-grade business flow automation, particularly in mission-critical contexts.
@@ 3.1 Tool-Policy Mapper
(1) Flight reservations can be canceled within 24 hours of booking, or if the airline canceled the flight. Otherwise, (2) basic economy or economy flights can be canceled only if travel insurance is bought […], and (3) explicit customer confirmation should be obtained prior to canceling a flight."

### v2

@@ 1 Introduction
Our approach enforces policy adherence through a two-phase process illustrated in Figure 1: (1) an offline buildtime step automatically maps policy fragments to the relevant tools and generates policy validation code; (2) during runtime these validators—called ToolGuards—are executed before agent’s tool invocation. If a planned action violates a policy, the agent is prompted to self-reflect and revise its plan before proceeding. Ultimately, the deployed ToolGuards will prevent the agent from taking an action violating a policy. This design ensures predictable and transparent policy enforcement with limited manual effort.
We evaluate the proposed approach on the $\tau$-bench Airlines domain (Yao et al., 2024) with a large set of (naturalistic, real-world) policies, 14 tools, and diverse customer-service scenarios. [^2] The other $\tau$-bench domain of Retail is smaller and less diverse; hence we limit the scope of this work to Airlines. We show that this ambitious and challenging task pushes the limits of contemporary LLMs, present encouraging preliminary results, and outline directions for future research and practical steps needed for this approach to mature toward enterprise adoption. Our code and data are available at https://github.com/IBM/tool_guard.
@@ 2 Related Work
Prior art in the area of constraint and policy adherence in agentic workflows is very sparse, with two major studies, to the best of our knowledge, explicitly addressing the integration of company policies within agentic flows (contrary to post-hoc remedy). Yao et al. (2024) were among the pioneers in this field, introducing a framework for evaluation of company policies adherence in multi-turn simulated flows in two domains: Airlines and Retail. Policies are represented by short documents, where domain-specific rules and constraints are detailed in natural language. Although the study reports failure rates stemming from a range of issues, the authors estimate that 25% of failures occur since the agent "fails to understand the domain-specific knowledge or rules". A recent work by Li et al. (2025) offers a framework for evaluation of policy adherence in multi-step single-turn tasks in five diverse domains, focusing on various composition types of policies ("single", "and", "or", "chain"). Here, concise and manually-crafted policies, as well as their compositions, are specified in configuration files. The authors report relatively low pass-rate across agents, varying from 31% to 69%. Both studies introduce policies to the agent in its prompt, instructing the LLM to follow the guidelines while satisfying user requirements.
Contrary to the pro-active policy enforcement above, another line of studies take the post-hoc reflection approach, by designing guardrails that check if any action taken by an agent violates a predefined policy (Huang et al., 2024; Rothkopf et al., 2024; Hua et al., 2024; Hoover et al., 2025). While this strategy is often easier for implementation (e.g., detecting that an unauthorized refund was issued or is simpler than preventing it), it may fall short of the strict demands of enterprise-grade business flow automation, particularly in mission-critical contexts, where a post-hoc response to a prohibited action turns often as "too late".
@@ 3.1 Tool-Policy Mapper
(1) Flight reservations can be canceled within 24 hours of booking, or if the airline canceled the flight. Otherwise, (2) basic economy or economy flights can be canceled only if travel insurance is bought […], and (3) explicit customer confirmation should be obtained prior to canceling a flight.
@@
Acknowledgements
We are thankful to Koren Lazar for his comments on an early version of this paper. We are also grateful to our three anonymous reviewers for their useful comments and constructive feedback.